Fermi National Laboratory
 

Accelerator Update for February 1, 2006 - February 3, 2006

 

Wednesday February 1

The day shift began with Operations monitoring the Antiproton Source (Pbar) stack and the Recycler stash, with the Tevatron (TeV) in shot setup, with Pbar experts investigating problems with the Lithium Lens, with Switchyard (SY) off due to shot setup, and with NuMI and MiniBooNE taking beam.

At 8:35 AM, experts accessed the Pbar AP0 target vault.

Operations established store 4618 at 10:55 AM with an initial luminosity of 110.6E30.

Operations resumed sending beam to the Meson MTest and MCenter experiments at 11:09 AM.

At 2:21 PM, a Booster RF power supply (BRF5) tripped off. Experts began making repairs. Booster beam resumed at 3:55 PM, but BRF5 remained off.

Earlier in the afternoon the NUMI enclosures were opened for a tour. At 4:11 PM, operators reported a problem with a door interlock. Safety system technicians investigated.

Pbar experts reported that the Lithium Lens repairs were completed by 5:02 PM.

NUMI resumed taking beam at 5:28 PM.

Pbar resumed stacking at 5:54 PM.

BRF5 tripped off at 7:08 PM. When it tripped, the station took down all the other Booster RF stations. Operators had Booster running within a few minutes, but called in RF experts to investigate anyway.

The Recycler began stashing antiprotons at 7:51 PM. The stash ended and Pbar resumed stacking at 8:45 PM.

At 9:51 PM, Operations reported that stacking had stopped. An expert reported that he thought the problem was due to a VCB. Experts investigated.

At 11:15 PM, store 4618 aborted. Operators investigated and found a bad CAMAC card.

At 11:36 PM, power supply experts had to switch out a feeder (MOS89) that halted all beam except for beam to MiniBooNE.

Thursday February 2

The midnight shift began with Operations monitoring the stack and stash, and with the F17B3 power supply work continuing. MiniBooNE was off at the beginning of the shift.

Operations prepared the D-Zero (D0) experiment at 12:01 AM for a collision hall access.

MiniBooNE resumed taking beam at 12:26 AM.

D0 came out of access at 2:03 AM.

Operations reported that power supply experts completed repairs on F17B3 at 4:50 AM.

Operations put the TeV into shot setup at 4:51 AM.

NuMI resumed taking beam at 5:26 AM.

At 7:34 AM, Pbar resumed stacking due to problems with the TeV shot. Investigations continue.

Operations put the TeV back into shot setup at 10:34 AM.

The NUMI experiment reported at 12:04 PM that they had finished their low intensity run. NUMI beam was turned off for an access.

Operations established store 4621 at 12:35 PM with an initial luminosity of 67.95E30.

The Recycler stashed antiprotons at 1:50 PM. The stash ended and Pbar began stacking at 2:59 PM. Beam to SY resumed two minutes later.

NUMI resumed taking beam at 3:26 PM.

At 4:48 PM, Operations reported problems with the Booster RF system. Experts investigated.

Operations turned off BRF4 and noted at 6:33 PM that the Booster RF problems cleared up.

The Recycler stashed antiprotons at 7:52 PM. The stash ended and Pbar began stacking at 8:47 PM.

At 10:56 PM, the Tevatron Electron Lens (TEL) tripped off. Operators investigated but found no bad status. An expert advised Operations to just keep reseting the device.

Friday February 3

The midnight shift began with Operations monitoring store 4621 and the stash, with Pbar stacking, and with MTest, MCenter, NuMI, and MiniBooNE taking beam.

The Recycler stashed antiprotons at 1:02 AM. The stash ended and Pbar resumed stacking at 1:56 AM. MTest and MCenter resumed taking beam.

The Recycler began preparing to stash antiprotons at 5:38 AM. The stash ended and Pbar resumed stacking around 6:30 AM. MTest and MCenter resumed taking beam.

TeV experts began an end-of-store study at 7:01 AM.

Operations terminated store 4621 at 8:15 AM.

The Plans for Friday and the Weekend

The plans are to establish a store later this morning after D0 completes their access. Pbar experts will be allowed to access their tunnel after the shot. The weekend will consist of stacking, stashing, and storing.

 
Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/03/2006    email Fermilab